OpenStudy (anonymous):

A. one solution B. infinite solutions C. no solution 3(4x-3)-7x=5x-9

OpenStudy (anonymous):

@bloonscat

OpenStudy (anonymous):

so if we simplify 3(4x-3)-7x=5x-9 by using the distributive property we get 12x - 9 - 7x = 5x - 9, and we can then simplify further by combining like terms, so 5x - 9 = 5x - 9. as we continue on, 5x - 5x = -9 + 9, which goes on to 0 = 0. Infinite solutions

OpenStudy (anonymous):

sure but isn't that against coc?

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

thk @bloonscat